OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the line is undefined, state so. (-2, 3) and (-4, 6)

OpenStudy (chaise):

m(slope) = y1-y2/x1-x2 m = 3-6/-2--4 m = :)

OpenStudy (anonymous):

Slope is defined at being the 2nd y-value minus the first divided by the 2nd x-value minus the first. With that said, we have the following equation: (6-3)/(-4-(-2)) This becomes the following fraction: 3/6 Which is 1/2 So your equation becomes: y = 1/2x + b, where b is the y-intercept. If this helps, please choose good answer. http://www.tutorsean.net